that u have to solve this problem by using differentiation and gradient at point on curves
1) At what point on the curve y=x^2+7x-8 is the gradient equal to 1 ?
2) At which points on the curve y=x^3+3x^2-5 is the gradient equal to 9 ?


Answer by Alan3354(69443)

1) At what point on the curve y=x^2+7x-8 is the gradient equal to 1 ?
y' = 2x + 7
2x + 7 = 1
x = -3
--> (-3,-20)
---------------------
2) At which points on the curve y=x^3+3x^2-5 is the gradient equal to 9 ?
y' = 3x^2 + 6x
3x^2 + 6x - 9 = 0
x^2 + 2x - 3 = 0
x = -3, 1
etc